# Multimodal Distributions for Circular Axial Data

### v1

@@ 3.2 Face Cleat in a Coal Seam
Face cleats are longitudinal fractures that represent significant geological features within coal seams-structures composed of coal deposits embedded within layers of rock. Variations in the orientation of these fractures can indicate hazardous mining conditions, including the potential presence of gas. Dataset B22 from Fisher (1993) includes 63 median directions of face cleats measured at the Wallsend Borehole Colliery in New South Wales, Australia. The median directions were measured at intervals of 20 meters along the coal seam. Consequently, the dataset exhibits sequential ordering, implying that the observations may not be independent. Despite the presence of potential dependence structures within the dataset, we analyzed the data as if it were a random sample to facilitate a direct comparison with results presented by Arnold and SenGupta (2006).
Table 1 (columns 8 to 10) provides maximized log-likelihood, Akaike Information Criterion (AIC), and Bayesian Information Criterion (BIC) values for fitted NNTS axial models with parameter $M$ ranging from 0 to 7. Figure 1 (right plot) shows the histogram along with the best-fitting AIC and BIC models, both of which select $M=6$. The selected NNTS model with $M=6$ displays two prominent modes within the observed range of data. This result contrasts with findings by Arnold and SenGupta (2006), who considered only unimodal densities for the same dataset.
@@ 3.3 Orientations of Termite Mounds
Dataset B13 from Fisher (1993) contains measurements of termite mound orientations for the species Amitermes laurensis, collected from 14 sites on the Cape York Peninsula, North Queensland, Australia. Observed orientations were standardized to the interval between 0 and $\pi$ by subtracting $\pi$ from any measurement greater than $\pi$. Figure 2 displays histograms for the orientation data at each of the 14 sites, along with sample sizes and geographical coordinates (latitude and longitude in degrees). The figure also shows the best-fitting NNTS axial densities according to the Bayesian Information Criterion (BIC) for each site. The final plot in Figure 2 presents the circular dot plot, histogram and the best BIC-fitted NNTS axial density for the combined data from all 14 sites, which serves as the basis for a likelihood ratio test of homogeneity, with the null hypothesis stating that the NNTS axial density is the same across all 14 sites. Table 2 provides maximized log-likelihood, AIC, and BIC values for NNTS axial densities with $M=0,1,\ldots,5$.
@@
Applying a likelihood ratio test for homogeneity using the best-fitting BIC-selected NNTS models for the 14 individual sites compared to the best-fitting BIC NNTS model for the combined dataset yields a p-value of $2.3648\times 10^{-8}$, assuming an asymptotic chi-squared distribution with 140 degrees of freedom. An alternative test, comparing the fitted NNTS models with $M=5$ at each site against a joint model, produces a p-value of 0.0062 with 130 degrees of freedom. In both tests, the null hypothesis of homogeneity across the 14 sites is rejected, indicating that variations exist in the orientation densities of termite mounds depending on their geographic location (latitude and longitude). These findings confirm the results obtained by Fisher (1993), who performed separate tests for the equality of mean directions and dispersions using a von Mises density model by doubling the axial angles to analyze them as circular data. However, it should be noted that the von Mises model is inherently unimodal and therefore unsuitable for modeling multimodal datasets, such as several of those shown in Figure 2, particularly evident in the histograms and fitted NNTS axial densities for site 8.
@@
Fisher (1993) conducted tests of homogeneity specifically for pairs of sites 5 and 14, and sites 6 and 8. Using a von Mises density model by doubling the axial angles to analyze them as circular data, Fisher rejected the null hypothesis of equal mean directions for both pairs. However, for the pair of sites 6 and 8, Fisher concluded that the hypothesis of equal dispersions could not be rejected. To comprehensively compare our results with Fisher’s (1993) analysis, we performed likelihood ratio homogeneity tests for the same pairs (sites 5 and 14, and sites 6 and 8) using an NNTS axial model with parameter $M=4$. The resulting p-values were 0.0051 for sites 5 and 14, and 0.0010 for sites 6 and 8, leading us to reject the null hypothesis of homogeneity for both pairs. These outcomes align with Fisher’s results, although Fisher’s tests separately evaluated mean directions and dispersions, while our approach employs a single integrated likelihood ratio test of overall homogeneity.
@@ 3.4 Magnetic Orientations of Ruminants
Table 3 shows maximized log-likelihood, AIC, and BIC values for fitted NNTS axial models with different values of $M$ (from 0 to 5 for red deer, and from 0 to 10 for cattle and roe deer). When testing homogeneity across the three populations, we clearly rejected the null hypothesis, obtaining p-values of $4.85\times 10^{-22}$ (best BIC model: $M=1$ for cattle, $M=3$ for red deer, $M=6$ for roe deer, and $M=4$ for all three populations combined) and $8.31\times 10^{-22}$ (best AIC model: $M=2$ for cattle, $M=6$ for red deer, $M=5$ for roe deer, and $M=5$ for all three populations combined).
When testing homogeneity specifically between red deer and roe deer, we obtained p-values of 0.0511 (best BIC model with $M=6$) and 0.9999 (best AIC model with $M=6$). Figure 3 displays circular dot plots, histograms, and fitted density curves based on the best BIC and AIC NNTS axial models for cattle, red deer, and roe deer. For each population, the modes correspond to the average magnetic declination of the various sampling locations worldwide, as reported by Begall et al. (2008).
@@ 3.5 Leaf Inclination Angles
The spatial orientation of plant leaves, particularly leaf inclination angles, provides essential information regarding key biological processes such as photosynthesis efficiency, reflectance, temperature regulation, and overall resource utilization. Leaf inclination angle distributions vary according to plant species and geographic location. Recent advancements in digital imaging technology have significantly increased the availability of leaf inclination angle measurements, as algorithms can now efficiently determine these angles directly from digital images (see Pisek and Adamson, 2020).
@@
In botanical studies, the leaf inclination angle $\theta_{leaf}$ typically ranges from 0 to $\frac{\pi}{2}$. It is common practice in the literature to model the transformed linear variable $X=\frac{2\theta_{leaf}}{\pi}$ using a Beta distribution with parameters $\alpha$ and $\beta$, and a probability density proportional to $x^{\alpha-1}(1-x)^{\beta-1}$. Based on parameter estimates for different plant species, de Wit (1965) proposed six theoretical distributions for leaf orientation angles: planophile (leaves predominantly horizontal), erectophile (leaves predominantly vertical), plagiophile (leaves predominantly inclined), extremophile (leaves predominantly horizontal and vertical), spherical (leaves oriented isotropically like a sphere), and uniform. The parameters of the Beta distribution for the plagiophile case satisfy $\alpha=\beta$ with $\alpha,\beta>1$, whereas for the extremophile case, they satisfy $\alpha=\beta$ with $\alpha,\beta<1$, indicating that the Beta distribution takes equal values at the extremes 0 and 1 (see Chianucci et al., 2018). For the uniform distribution, the parameters satisfy $\alpha=\beta=1$. Therefore, the angle $\theta=2\theta_{leaf}$ can be treated as an axial angle, and the distributions of $\theta$ for plants exhibiting plagiophile, extremophile, or uniform leaf distributions can be modeled using NNTS axial models.
@@
The dataset from Pisek and Adamson (2020) contains leaf inclination angle measurements for 71 species of gum trees (genus Eucalyptus). From these, we selected 11 species observed at the Huntington Library, Art Collections, and Botanical Gardens in Pasadena, California (latitude 34.125, longitude -118.114, altitude 207 m.a.s.l.), classified by Pisek and Adamson (2020) as either uniform (9 species) or plagiophile (2 species). Table 4 provides species names, sample sizes, Beta distribution parameter estimates, the optimal $M$ values for the best-fitting NNTS axial models based on the Bayesian Information Criterion (BIC), and the corresponding BIC values. Figures 4 and 5 display circular dot plots, histograms, and the best BIC-fitted NNTS axial densities for species classified as uniform (U) and plagiophile (PG), respectively.
@@
Notably, for species classified as uniform (cases 1 balladoniensis and 10 shirleyi), neither the best BIC- nor AIC-selected NNTS axial densities corresponded to a uniform distribution (Figure 4). This observation is further confirmed by the uniformity test of Fernández-Durán and Gregorio-Domínguez (2024) shown in Table 4, where the null hypothesis of uniformity is rejected for cases 1 and 10 when comparing against either a NNTS axial model with $M=1$ or the best BIC-selected NNTS axial model. Moreover, when using the best AIC-selected NNTS axial model as an alternative, the hypothesis of uniformity is additionally rejected for cases 2, 6, and 8 at a 5% significance level.
@@ 4 Conclusions
A flexible new family of axial data densities is developed by identifying conditions under which NNTS densities, originally defined for circular data, are suitable for axial data. These NNTS axial densities may be uniform, symmetric, asymmetric, or multimodal, depending on parameter constraints. Efficient numerical optimization on manifolds, adapted from Fernández-Durán and Gregorio-Domínguez (2010), facilitates maximum likelihood estimation and enables direct implementation of likelihood ratio tests. The proposed methodology was successfully applied to real datasets involving the orientations of rocks, animals, and plants, showing excellent performance. Importantly, these results demonstrate significant improvements over previous methodologies that did not consider multimodal axial densities.

### v2

@@ 3.2 Face Cleat in a Coal Seam
Face cleats are longitudinal fractures that represent significant geological features within coal seams-structures composed of coal deposits embedded within layers of rock. Variations in the orientation of these fractures can indicate hazardous mining conditions, including the potential presence of gas. Dataset B22 from Fisher (1993) includes 63 median directions of face cleats measured at the Wallsend Borehole Colliery in New South Wales, Australia. The median directions were measured at intervals of 20 meters along the coal seam. Consequently, the dataset exhibits sequential ordering, implying that the observations may not be independent. Despite the presence of potential dependence structures within the dataset, we analyzed the data as if it were a random sample to facilitate a direct comparison with results presented by Arnold & SenGupta (2006).
Table 1 (columns 8 to 10) provides maximized log-likelihood, Akaike Information Criterion (AIC), and Bayesian Information Criterion (BIC) values for fitted NNTS axial models with parameter $M$ ranging from 0 to 7. Figure 1 (right plot) shows the histogram along with the best-fitting AIC and BIC models, both of which select $M=6$. The selected NNTS model with $M=6$ displays two prominent modes within the observed range of data. This result contrasts with findings by Arnold & SenGupta (2006), who considered only unimodal densities for the same dataset.
@@ 3.3 Orientations of Termite Mounds
Dataset B13 from Fisher (1993) contains measurements of termite mound orientations for the species Amitermes laurensis, collected from 14 sites on the Cape York Peninsula, North Queensland, Australia. Observed orientations were standardized to the interval between 0 and $\pi$ by subtracting $\pi$ from any measurement greater than $\pi$. Figure 2 displays histograms for the orientation data at each of the 14 sites, along with sample sizes and geographical coordinates (latitude and longitude in degrees). The figure also shows the best-fitting NNTS axial densities according to the Bayesian Information Criterion (BIC) for each site. The final plot in Figure 2 presents the histogram and the best BIC-fitted NNTS axial density for the combined data from all 14 sites, which serves as the basis for a likelihood ratio test of homogeneity, with the null hypothesis stating that the NNTS axial density is the same across all 14 sites. Table 3 provides maximized log-likelihood, AIC, and BIC values for NNTS axial densities with $M=0,1,\ldots,8$.
Table 4 (column 5) reports the $p$-values from the likelihood ratio tests of homogeneity based on NNTS axial models for the 14 individual sites, with $M_{\text{pooled}}$ ranging from 1 to 8. For example, comparing the fitted NNTS models with $M=5$ at each site against a joint model with $M_{\text{pooled}}=5$ yields a $p$-value of 0.0062 with 130 degrees of freedom (see column 10 in Table 3). For the cases where $M_{k}=M_{\text{pooled}}=4,5,6,$ and $7$ for $k=1,2,\ldots,14$, the null hypothesis of homogeneity across the 14 sites is rejected at the 5% significance level, indicating that the orientation densities of termite mounds vary with geographic location (latitude and longitude). The multiple-testing-corrected $p$-values, considering the eight tests, are 0.0400 under the Bonferroni correction and 0.0393 under the Tippett correction, both leading to rejection of the null hypothesis of homogeneity at the 5% significance level.
@@
These findings confirm the results obtained by Fisher (1993), who performed separate tests for the equality of mean directions and dispersions using a von Mises density model by doubling the axial angles to analyze them as circular data. However, it should be noted that the von Mises model is inherently unimodal and therefore unsuitable for modeling multimodal datasets, such as several of those shown in Figure 2, particularly evident in the histograms and fitted NNTS axial densities for site 8.
@@
Fisher (1993) conducted homogeneity tests specifically for the pairs of sites 5 and 14, and sites 6 and 8. Using a von Mises density model and doubling the axial angles to treat them as circular data, Fisher rejected the null hypothesis of equal mean directions for both pairs. However, for sites 6 and 8, Fisher (1993) concluded that the hypothesis of equal dispersions could not be rejected. To enable a direct comparison with Fisher’s (1993) analysis, we performed NNTS axial likelihood ratio homogeneity tests for the same site pairs (5 and 14, and 6 and 8), using NNTS axial models with $M_{\text{pooled}}$ ranging from 1 to 8. The resulting $p$-values, shown in Table 4 (columns 9 and 13), were all greater than 0.05, with a minimum of 0.4428 for sites 5 and 14 and a minimum of 0.7895 for sites 6 and 8. Consequently, the null hypothesis of homogeneity was not rejected for either pair, even when considering the Bonferroni- and Tippett-corrected $p$-values. These findings differ from those reported by Fisher (1993), as Fisher’s tests separately assessed equality of mean directions and dispersions, whereas our approach applies a unified likelihood ratio test for overall homogeneity.
@@ 3.4 Magnetic Orientations of Ruminants
Table 5 shows maximized log-likelihood, AIC, and BIC values for fitted NNTS axial models with different values of $M$ (from 0 to 6 for red deer, and from 0 to 10 for cattle and roe deer).
For testing homogeneity, Table 6 presents the maximized log-likelihood, BIC, and AIC values for the NNTS models fitted to the pooled populations of cattle, red deer, and roe deer (columns 2–5), and to the pooled populations of red and roe deer (columns 6–9), for $M_{\text{pooled}}$ ranging from 1 to 10. For the three-population case (cattle, red deer, and roe deer), all individual $p$-values (column 5) are well below 0.05, rejecting the null hypothesis of homogeneity, except for $M_{\text{pooled}}=1$, which yields a $p$-value of 0.0731. In contrast, for the two-population case (red and roe deer), all individual $p$-values (column 9) are well above 0.05, indicating that the null hypothesis of homogeneity cannot be rejected at the 5% significance level. For the three-population comparison, the Bonferroni- and Tippett-corrected $p$-values are $2.13\times 10^{-19}$ and <0.0001<0.0001, respectively. For the two-population comparison (red and roe deer), the corresponding corrected $p$-values are 1.0000 and 0.9999, respectively.
Figure 3 displays circular dot plots, histograms, and fitted density curves based on the best BIC and AIC NNTS axial models for cattle, red deer, and roe deer. For each population, the modes correspond to the average magnetic declination of the various sampling locations worldwide, as reported by Begall et al. (2008).
@@ 3.5 Leaf Inclination Angles
The spatial orientation of plant leaves, particularly leaf inclination angles, provides essential information regarding key biological processes such as photosynthesis efficiency, reflectance, temperature regulation, and overall resource utilization. Leaf inclination angle distributions vary according to plant species and geographic location. Recent advancements in digital imaging technology have significantly increased the availability of leaf inclination angle measurements, as algorithms can now efficiently determine these angles directly from digital images (see Pisek & Adamson, 2020).
@@
In botanical studies, the leaf inclination angle $\theta_{leaf}$ typically ranges from 0 to $\frac{\pi}{2}$. It is common practice in the literature to model the transformed linear variable $X=\frac{2\theta_{leaf}}{\pi}$ using a Beta distribution with parameters $\alpha$ and $\beta$, and a probability density proportional to $x^{\alpha-1}(1-x)^{\beta-1}$. Based on parameter estimates for different plant species, de Wit (1965) proposed six theoretical distributions for leaf orientation angles: planophile (leaves predominantly horizontal), erectophile (leaves predominantly vertical), plagiophile (leaves predominantly inclined), extremophile (leaves predominantly horizontal and vertical), spherical (leaves oriented isotropically like a sphere), and uniform. The parameters of the Beta distribution for the plagiophile case satisfy $\alpha=\beta$ with $\alpha,\beta>1$, whereas for the extremophile case, they satisfy $\alpha=\beta$ with $\alpha,\beta<1$, indicating that the Beta distribution takes equal values at the extremes 0 and 1 (see Chianucci et al., 2018). For the uniform distribution, the parameters satisfy $\alpha=\beta=1$. Therefore, for plants exhibiting plagiophile, extremophile, or uniform leaf distributions, the transformed angle $\theta=2\theta_{leaf}$ can be treated as an axial angle, given the $\pi$-periodicity of its density function, and the NNTS axial models can be applied to model the distribution of $\theta$.
@@
The dataset from Pisek & Adamson (2020) contains leaf inclination angle measurements for 71 species of gum trees (genus Eucalyptus). From these, we selected 11 species observed at the Huntington Library, Art Collections, and Botanical Gardens in Pasadena, California (latitude 34.125, longitude -118.114, altitude 207 m.a.s.l.), classified by Pisek & Adamson (2020) as either uniform (9 species) or plagiophile (2 species). Table 7 provides species names, sample sizes, Beta distribution parameter estimates, the optimal $M$ values for the best-fitting NNTS axial models based on the Bayesian Information Criterion (BIC), and the corresponding BIC values. Figures 4 and 5 display circular dot plots, histograms, and the best BIC-fitted NNTS axial densities for species classified as uniform (U) and plagiophile (PG), respectively.
@@
Notably, for species classified as uniform (cases 1 balladoniensis and 10 shirleyi), neither the best BIC- nor AIC-selected NNTS axial densities corresponded to a uniform distribution (Figure 4). This observation is further confirmed by the uniformity test of Fernández-Durán & Gregorio-Domínguez (2024) shown in Table 7, where the null hypothesis of uniformity is rejected for cases 1 and 10 when comparing against either a NNTS axial model with $M=1$ or the best BIC-selected NNTS axial model. Moreover, when using the best AIC-selected NNTS axial model as an alternative, the hypothesis of uniformity is additionally rejected for cases 2, 6, and 8 at a 5% significance level.
Table 8 presents the $p$-values of the homogeneity tests for the nine cases of U species (column 5), the two cases of PG species (column 9), and the combined eleven cases of U and PG species (column 13), for values of the parameter $M_{\text{pooled}}$ ranging from 1 to 8. For the U species, the null hypothesis of homogeneity is clearly rejected, with Bonferroni- and Tippett-corrected $p$-values both equal to 2.6028e-07. For the PG species, the individual tests for $M_{\text{pooled}}=1,\ldots,8$ do not reject the null hypothesis of homogeneity. The Bonferroni- and Tippett-corrected $p$-values are equal to 1.0000 and 0.7924, respectively, indicating no rejection of the null hypothesis of homogeneity. Finally, for the combined U and PG species, as expected, both the individual and corrected $p$-values are very small, clearly rejecting the null hypothesis of homogeneity.
@@ 4 Conclusions
A flexible new family of axial data densities is developed by identifying conditions under which NNTS densities, originally defined for circular data, are suitable for axial data. These NNTS axial densities may be uniform, symmetric, asymmetric, or multimodal, depending on parameter constraints. Efficient numerical optimization on manifolds, adapted from Fernández-Durán & Gregorio-Domínguez (2010), facilitates maximum likelihood estimation and enables direct implementation of likelihood ratio tests. The proposed methodology was successfully applied to real datasets involving the orientations of rocks, animals, and plants, showing excellent performance. Importantly, these results demonstrate significant improvements over previous methodologies that did not consider multimodal axial densities.